THE INFRARED SPECTRUM OF HBO

\underline{P. Colarusso}, K.-Q. Zhang, and P.F. Bernath, Centre for Molecular Beams and Laser Chemistry, Univeristy of Waterloo, Waterloo, Ontario, Canada, N2L 3G1.

The free radical HBO was generated in a high-temperature furnace at 1400^{∘}C. The infrared spectra of the two major isotopic species, H^{11}BO and H^{10}BO, were recorded in emission with a Fourier transform spectrometer (Bruker IFS 120HR). The wavenumber range from 400 to 3500 cm^{−1} was detected at a resolution of 0.01 cm^{−1}. The three fundamental vibrational bands as well as various combination and hot bands have been observed. A vibrational-rotational analysis of the spectra will be presented.